OpenStudy (anonymous):

Need help with this problem that needs to be solved using substitution: 3/4x-1/12y=-5/2 3/8x-1/6y=-1/2 I came up with several different answers and really don't know what to do...I tried multiplying it with 12 or 24 and it just doesn't makes sense..please help, thanks.....

OpenStudy (anonymous):

fractions are fractious so get rid of them x ist by 12 9x - y = - 30............a x 2nd by 24 9x - 4y = -12..............b from a : 9x = y-30 plug this into b: y - 30 - 4y = -12 3y = -18 y = -6 now plug this into equation b to find x and you are home!

OpenStudy (anonymous):

Thanks for your help, I think I got it now.